# MAESTRO: Multi-modal Adaptive Ensemble for Spectro-Temporal Robust Optimization

### v1

@@ 3.9. Uncertainty Estimation
For point forecasts, the model is trained by minimizing a standard regression loss function between the predicted mean $\hat{y}_{t}$ and the true value $y_{t}$. We use either the Mean Squared Error (MSE) loss or the Huber loss.
@@
To incorporate uncertainty estimation, we modify the final prediction layer. Instead of a single output, the model employs a dedicated uncertainty head that, given the final feature representation $h_{t}\in\mathbb{R}^{d_{\text{model}}}$ for time step $t$, produces both a mean prediction $\hat{y}_{t}$ and a standard deviation $\sigma_{t}$. This is achieved through two separate linear projections:
@@
3.10. Error Analysis
To further dissect the performance of MAESTRO against the baseline models, we conducted a detailed error analysis, as depicted in Figure 2. The analysis includes a histogram of absolute errors, a box plot summarizing the error distributions, a cumulative distribution function (CDF) of the errors, and a statistical significance matrix comparing the models.
Figure 2. Model Prediction Error Distribution Comparison. This figure provides a comprehensive comparison of the absolute errors of MAESTRO and three baseline models (TimesNet, iTransformer, and DLinear). The top-left panel shows a histogram of the error distributions. The top-right panel presents a box plot for each model, highlighting the median, quartiles, and outliers. The bottom-right panel displays the cumulative distribution function (CDF) of the absolute errors, indicating the probability of the error being less than or equal to a certain value. The bottom-left panel shows the results of statistical significance tests (p-values) and effect sizes between the models.
4. Principled Architectural Design
In this section, we provide a theoretical rationale for MAESTRO’s architecture. We argue that its composition of spectro-temporal modules is not an arbitrary ensemble but a principled design tailored to the challenges of epidemiological forecasting: multi-scale seasonality, noisy data, and long-range dependencies. Our design philosophy is to assign specific, complementary roles to each component, guided by their theoretical properties.
@@ 4.1. Multi-Scale Seasonality Decomposition
Methodology: We first use a moving-average filter, $\mathcal{D}_{K}$, to detrend the series(Box et al., 2015; Wu et al., 2021), separating it into a low-frequency trend component $x^{(\mathrm{tr})}$ and a high-frequency detrended component $x^{(\mathrm{se})}$ that captures seasonal and residual variations:
(24) $x^{(\mathrm{tr})}$ $=\mathcal{D}_{K}(x)=\text{AvgPool1d}(\text{Padding}(x),\text{kernel\_size}=K,\text{stride}=1),\quad x^{(\mathrm{se})}=x-x^{(\mathrm{tr})}$
@@ 4.2. Hybrid Modeling of Long-Range and Aperiodic Dynamics
Methodology: The core of our design is a synergistic combination of Mamba blocks and a multi-scale Transformer encoder. The Mamba block is based on a continuous-time State-Space Model (SSM) defined by:
@@ 4.3. Expressiveness of State-Space and Decomposition
Following the successful application of decomposition in long-term forecasting (Wu et al., 2021; Zhou et al., 2022), we also simplify the learning task by initially decomposing the input $x$ into a trend component $x^{(\mathrm{tr})}=\mathcal{D}_{K}(x)$ and a seasonal component $x^{(\mathrm{se})}=x-x^{(\mathrm{tr})}$. The trend, as a low-frequency signal, can be effectively modeled by simpler network components. This allows the high-capacity encoders to focus on the more complex, high-frequency seasonal variations. Such separation can be viewed as a form of curriculum learning, which can improve model stability and sample efficiency.
@@ 4.4. Uncertainty Estimation
We train with a composite objective $\mathcal{L}=\lambda_{1}\,\operatorname{MSE}+\lambda_{2}\,\operatorname{NLL}+\lambda_{3}\,R_{\text{spectral}}+\lambda_{4}\,R_{\text{weights}}+\lambda_{5}\,R_{\text{stability}}$, where $R_{\text{spectral}}$ includes magnitude and smoothness penalties on $M(\omega)$; $R_{\text{weights}}$ regularizes ensemble weights (entropy/smoothness); and $R_{\text{stability}}$ constraints SSM spectra.
@@ 5. Experimental Setup
Classical: ARIMA; ML: SVR; RNN: LSTM, GRU; Transformers: Transformer, Informer, Autoformer, FEDformer, iTransformer; Spectro/Conv: TimesNet; Linear head: DLinear. We use official or widely used implementations with consistent preprocessing and hyperparameter budgets.
@@
5.1. Data Overview
To provide a comprehensive overview of the datasets utilized in this study, we present a series of visualizations that encapsulate their primary characteristics. These visualizations are crucial for understanding the underlying data distributions and temporal patterns that our proposed model, MAESTRO, aims to learn.
The visualizations in Figure 3 and Figure 4 offer a foundational understanding of the data’s structure. Figure 3 provides a macro-level view of the temporal trends, clearly revealing seasonal and residual components that justify our model’s decomposition-based architecture. Concurrently, Figure 4 details the statistical distributions, highlighting the pronounced right-skew and high kurtosis in ILI and Google Trends data, which underscores the need for robust loss functions and uncertainty estimation. These insights are instrumental in appreciating the complexity of the forecasting task and the design of our model.
5.2. Evaluation Methodology
We adopt chronological splits for the dataset, partitioning it into training (60%), validation (20%), and testing (20%) sets. We report the mean over 5 seeds. For fairness, we align look-back $L$, horizon $H$, optimizer, epochs, and early stopping across models. Modalities are concatenated for single-stream baselines; models with native multi-branch fusion are also evaluated in that mode when feasible.
@@ 6.1. Main Results
The primary results of our forecasting experiments are presented in Figure 5. This figure provides a comprehensive visualization of MAESTRO’s predictive performance against the ground truth, alongside several state-of-the-art baseline models. The visualization highlights the model’s ability to accurately capture the complex temporal patterns and fluctuations present in the influenza-like illness (ILI) data.
@@ 6.2. Performance Comparison
To provide a clear and quantitative comparison of MAESTRO’s performance against baseline models, we present the key performance metrics in Table 1, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), and the coefficient of determination (R2).
@@ 6.3. Ablation Studies
We follow consistent naming between paper and code for ablations:
@@
Table 2. Ablation study results. We compare MAESTRO (Full Model) against variants with specific components removed or isolated. Performance metrics (MAE, RMSE, R2) and parameter counts are reported. Lower is better for MAE/RMSE, higher for R2.
@@
Figure 6. Comprehensive prediction comparison across ablations and baselines. The red dashed line separates training and testing periods; R2 is annotated in each panel.
[…]
Multi-panel figure showing original, fitted, and predicted influenza time series across MAESTRO variants and baselines, with a red dashed split line and per-panel R-squared labels.
@@ 7. Discussion and Limitations
Our results highlight the trade-offs inherent in multi-metric evaluation. While MAESTRO achieves a state-of-the-art R2 (0.956), indicating a superior overall fit to the data’s variance, it is marginally outperformed by TimesNet on MAE and RMSE. This suggests that while our spectro-temporal fusion architecture excels at capturing the overall trend and seasonality, other models may be more specialized in minimizing point-wise absolute errors. The universally high MAPE values, as seen in Table 1, are a known artifact of time series with near-zero values (common in off-seasons), which can disproportionately inflate the metric. We thus contend that R2, MAE, and RMSE are more reliable indicators for this forecasting task, with R2 underscoring our model’s strength in explaining the underlying data structure.
The architectural gains of MAESTRO are primarily demonstrated empirically through extensive ablation studies (see Figure 7). We provided a preliminary analysis of spectro-temporal factorization, efficient long-range modeling via SSM, and cross-modal fusion. However, a formal theoretical analysis of why this specific combination of components is effective, particularly regarding its generalization performance under distribution shifts, remains a promising future research direction. The model’s complexity, while justified by its performance, also warrants further investigation into more lightweight yet equally robust configurations.
Our study is conducted on data from a single geography (Hong Kong), which limits the direct generalizability of our findings to other regions with different climates, healthcare systems, or population behaviors. Cross-region validation and domain adaptation are therefore critical next steps. Furthermore, while we intentionally excluded the COVID-19 period to ensure a stable training regime, the model’s robustness to unforeseen structural breaks and regime shifts, such as pandemics, has not been fully evaluated and merits dedicated future study. This naturally leads to research on adaptive learning horizons and online model retraining strategies.
8. Ethical Considerations
This work uses aggregated, publicly available surveillance, search trends, and meteorological data. Predictions should complement, not replace, expert epidemiological judgment. We caution against over-reliance in settings with substantial reporting delays or structural breaks.
@@ 9. Conclusion
We introduced MAESTRO, a spectro-temporal multi-modal framework that fuses surveillance, search trends, and weather for influenza forecasting. The architecture couples decomposition, attention, state-space modeling, multi-scale convolutions, and frequency-domain learning. On Hong Kong data, it achieves a state-of-the-art R2 of 0.956 while demonstrating competitive performance on other metrics. Future work includes cross-region transfer, adaptive horizons, and theoretical analysis of spectro-temporal fusion.
10. Acknowledgments
We thank the Hong Kong Department of Health for influenza surveillance, Google Trends for search data, and the Hong Kong Observatory for meteorological data. This work was supported in part by institutional funds from Macau University of Science and Technology. We also thank colleagues for helpful discussions.

### v3

@@ 3.9. Uncertainty Estimation
For point forecasts, the model is trained by minimizing a standard regression loss function between the predicted mean $\hat{y}_{t}$ and the true value $y_{t}$. Either the Mean Squared Error (MSE) loss or the Huber loss is used.
@@
To incorporate uncertainty estimation, the final prediction layer is modified. Instead of a single output, the model employs a dedicated uncertainty head that, given the final feature representation $h_{t}\in\mathbb{R}^{d_{\text{model}}}$ for time step $t$, produces both a mean prediction $\hat{y}_{t}$ and a standard deviation $\sigma_{t}$. This is achieved through two separate linear projections:
@@
4. Architectural Design Principles
In this section, a theoretical rationale is provided for MAESTRO’s architecture. The argument is that its composition of spectro-temporal modules is not an arbitrary ensemble but a principled design tailored to the challenges of epidemiological forecasting: multi-scale seasonality, noisy data, and long-range dependencies. The design philosophy is to assign specific, complementary roles to each component, guided by their theoretical properties.
@@ 4.1. Multi-Scale Seasonality Decomposition
Methodology: The model employs an additive decomposition strategy, inspired by classical methods like STL (Seasonal-Trend decomposition based on Loess) (Cleveland et al., 1990). Both approaches share the foundational goal of separating a time series $x_{t}$ into trend and seasonal components. However, they differ significantly in their methodology for achieving this separation.
A classical STL decomposition follows an iterative procedure to robustly separate a series $x_{t}$ into a trend $T_{t}$, a seasonal $S_{t}$, and a remainder component $R_{t}$, such that $x_{t}=T_{t}+S_{t}+R_{t}$. The core of STL lies in its use of Loess (Locally Estimated Scatterplot Smoothing), a non-parametric regression method. The process involves an inner loop for seasonal and trend smoothing and an outer loop that computes robustness weights to mitigate the effect of outliers. Formally, at each outer loop iteration $j$, the trend and seasonal components are updated via a sequence of Loess smoothing operations, denoted by the operator $\mathcal{L}$:
(24) $S_{t}^{(j)}$ $=\mathcal{L}_{\text{seasonal}}(x_{t}-T_{t}^{(j-1)})$
(25) $T_{t}^{(j)}$ $=\mathcal{L}_{\text{trend}}(x_{t}-S_{t}^{(j)})$
This iterative, weighted regression makes STL robust but computationally intensive.
In contrast, MAESTRO adopts the decomposition principle but replaces the iterative Loess smoother with a single-pass, linear moving-average filter for computational efficiency within a deep learning framework. This filter, $\mathcal{D}_{K}$, is implemented as a 1D average pooling layer. The decomposition is direct and non-iterative:
(26) $x^{(\mathrm{tr})}_{t}$ $=\mathcal{D}_{K}(x_{t})=\text{AvgPool1d}(\text{Padding}(x_{t}),\text{kernel\_size}=K,\text{stride}=1)$
(27) $x^{(\mathrm{se})}_{t}$ $=x_{t}-x^{(\mathrm{tr})}_{t}$
While this approach is less robust to outliers than STL, its computational efficiency and differentiability make it highly suitable for end-to-end training of large-scale models. It effectively captures the low-frequency trend, aligning with the primary goal of decomposition in forecasting models (Wu et al., 2021; Zeng et al., 2023).
@@ 4.2. Hybrid Modeling of Long-Range and Aperiodic Dynamics
Methodology: The core of the design is a synergistic combination of Mamba blocks and a multi-scale Transformer encoder. The Mamba block is based on a continuous-time State-Space Model (SSM) defined by:
@@ 4.3. Expressiveness of State-Space and Decomposition
Following the successful application of decomposition in long-term forecasting (Wu et al., 2021; Zhou et al., 2022), the learning task is also simplified by initially decomposing the input $x$ into a trend component $x^{(\mathrm{tr})}=\mathcal{D}_{K}(x)$ and a seasonal component $x^{(\mathrm{se})}=x-x^{(\mathrm{tr})}$. The trend, as a low-frequency signal, can be effectively modeled by simpler network components. This allows the high-capacity encoders to focus on the more complex, high-frequency seasonal variations. Such separation can be viewed as a form of curriculum learning, which can improve model stability and sample efficiency.
@@ 4.4. Uncertainty Estimation
The model is trained with a composite objective $\mathcal{L}=\lambda_{1}\,\operatorname{MSE}+\lambda_{2}\,\operatorname{NLL}+\lambda_{3}\,R_{\text{spectral}}+\lambda_{4}\,R_{\text{weights}}+\lambda_{5}\,R_{\text{stability}}$, where $R_{\text{spectral}}$ includes magnitude and smoothness penalties on $M(\omega)$; $R_{\text{weights}}$ regularizes ensemble weights (entropy/smoothness); and $R_{\text{stability}}$ constraints SSM spectra.
@@
5. Data Overview
To provide a comprehensive overview of the datasets utilized in this study, a series of visualizations are presented that encapsulate their primary characteristics. These visualizations are crucial for understanding the underlying data distributions and temporal patterns that the proposed model, MAESTRO, aims to learn.
The visualizations in Figure 2 and Figure 3 offer a foundational understanding of the data’s structure. Figure 2 provides a macro-level view of the temporal trends, clearly revealing seasonal and residual components that justify the model’s decomposition-based architecture. Concurrently, Figure 3 details the statistical distributions, highlighting the pronounced right-skew and high kurtosis in ILI and Google Trends data, which underscores the need for robust loss functions and uncertainty estimation. These insights are instrumental in appreciating the complexity of the forecasting task and the design of the model.
@@ 6. Experimental Setup
Chronological splits are adopted for the dataset, partitioning it into training (60%), validation (20%), and testing (20%) sets. The mean performance over 5 independent training runs is reported, each with a different random seed, to ensure the stability of the results. For fairness, look-back $L$, horizon $H$, optimizer, epochs, and early stopping are aligned across models. Modalities are concatenated for single-stream baselines; models with native multi-branch fusion are also evaluated in that mode when feasible.
@@
Classical: ARIMA; ML: SVR; RNN: LSTM, GRU; Transformers: Transformer, Informer, Autoformer, FEDformer, iTransformer; Spectro/Conv: TimesNet; Linear head: DLinear. Official or widely used implementations are used with consistent preprocessing and hyperparameter budgets.
@@ 7.1. Main Results
The primary results of the forecasting experiments are presented in Figure 4. This figure provides a comprehensive visualization of MAESTRO’s predictive performance against the ground truth, alongside several state-of-the-art baseline models. The visualization highlights the model’s ability to accurately capture the complex temporal patterns and fluctuations present in the influenza-like illness (ILI) data.
@@ 7.2. Performance Comparison
To provide a clear and quantitative comparison of MAESTRO’s performance against baseline models, the key performance metrics are presented in Table 1, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), and the coefficient of determination (R2).
@@ 7.3. Ablation Studies
Consistent naming is followed between paper and code for ablations:
@@
Table 2. Ablation study results. MAESTRO (Full Model) is compared against variants with specific components removed or isolated. Performance metrics (MAE, RMSE, R2) and parameter counts are reported. Lower is better for MAE/RMSE, higher for R2.
@@
Figure 5. Comprehensive prediction comparison across ablations and baselines. The red dashed line separates training and testing periods; R2 is annotated in each panel.:
@@
7.4. Error Analysis
To further dissect the performance of MAESTRO against the baseline models, a detailed error analysis was conducted, as depicted in Figure 7. The analysis includes a histogram of absolute errors, a box plot summarizing the error distributions, a cumulative distribution function (CDF) of the errors, and a statistical significance matrix comparing the models.
Figure 7. Model Prediction Error Distribution Comparison. This figure provides a comprehensive comparison of the absolute errors of MAESTRO and three baseline models (TimesNet, iTransformer, and DLinear). (A) A histogram of the error distributions. (B) A box plot for each model, highlighting the median, quartiles, and outliers. (C) The results of statistical significance tests (p-values and effect sizes between the models). (D) The cumulative distribution function (CDF) of the absolute errors, indicating the probability of the error being less than or equal to a certain value.
@@ 8. Discussion and Limitations
The results highlight the trade-offs inherent in multi-metric evaluation. While MAESTRO achieves a state-of-the-art R2 (0.956), indicating a superior overall fit to the data’s variance, it is marginally outperformed by TimesNet on MAE and RMSE. This suggests that while the spectro-temporal fusion architecture excels at capturing the overall trend and seasonality, other models may be more specialized in minimizing point-wise absolute errors. The universally high MAPE values, as seen in Table 1, are a known artifact of time series with near-zero values (common in off-seasons), which can disproportionately inflate the metric. For clarity, MAPE is reported calculated using the standard formula without any smoothing or adjustments for these near-zero values. It is thus contended that R2, MAE, and RMSE are more reliable indicators for this forecasting task, with R2 underscoring the model’s strength in explaining the underlying data structure.
The architectural gains of MAESTRO are primarily demonstrated empirically through extensive ablation studies (see Figure 6). A preliminary analysis of spectro-temporal factorization, efficient long-range modeling via SSM, and cross-modal fusion was provided. However, a formal theoretical analysis of why this specific combination of components is effective, particularly regarding its generalization performance under distribution shifts, remains a promising future research direction. The model’s complexity, while justified by its performance, also warrants further investigation into more lightweight yet equally robust configurations.
This study is conducted on data from a single geography (Hong Kong), which limits the direct generalizability of the findings to other regions with different climates, healthcare systems, or population behaviors. Cross-region validation and domain adaptation are therefore critical next steps. Furthermore, while the COVID-19 period was intentionally excluded to ensure a stable training regime, the model’s robustness to unforeseen structural breaks and regime shifts, such as pandemics, has not been fully evaluated and merits dedicated future study. This naturally leads to research on adaptive learning horizons and online model retraining strategies.
@@ 9. Conclusion
MAESTRO is introduced, a spectro-temporal multi-modal framework that fuses surveillance, search trends, and weather for influenza forecasting. The architecture couples decomposition, attention, state-space modeling, multi-scale convolutions, and frequency-domain learning. On Hong Kong data, it achieves a state-of-the-art R2 of 0.956 while demonstrating competitive performance on other metrics. Future work includes cross-region transfer, adaptive horizons, and theoretical analysis of spectro-temporal fusion.
Acknowledgements.
The Hong Kong Department of Health is acknowledged for influenza surveillance, Google Trends for search data, and the Hong Kong Observatory for meteorological data. This work was supported in part by institutional funds from Macau University of Science and Technology. Colleagues are also thanked for helpful discussions.
10. Online Resources
The source code for MAESTRO is publicly available in a GitHub repository: https://github.com/Hylouis233/MAESTRO.